OpenStudy (anonymous):

Statistics Assume that the population of heights of male college students is approximately normally distributed with mean m of 70 inches and standard deviation s of 3.65 inches. Show all work. (A) I have worked out. (B) Find the proportion of male college students whose height is no more than 68 inches.

OpenStudy (amistre64):

i found this one

OpenStudy (amistre64):

no more than 68 means 68 or less P(X<68)

OpenStudy (amistre64):

since the mean is at 70 this is to the left of the mean and will give us a negative zscore

OpenStudy (amistre64):

as before: z = (x-mean)/sd z = (68-70)/3.65 = -0.615

OpenStudy (anonymous):

ok, I couldn't find any examples in my book that said less than.

OpenStudy (amistre64):

from our last problem, you need to recall that when the z score is negative, we will be subtracting from .5 and not adding to it.

OpenStudy (amistre64):

what value do you get for when z = 0.615?

OpenStudy (amistre64):

i got no idea how to round that ...

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=zscore+%2868-70%29%2F3.25 wolfram gives some nice results

OpenStudy (anonymous):

just a minute let me look

OpenStudy (anonymous):

when i calculated it i got -2/3.65=-0.5479

OpenStudy (amistre64):

yeah, i typoed it into 3.25

OpenStudy (amistre64):

yours is right

OpenStudy (anonymous):

ok let me see here then: 0.2054

OpenStudy (amistre64):

id round up to .55 instead of down to .54

OpenStudy (anonymous):

oh yeah, i need to round to the nearest 0.00

OpenStudy (anonymous):

0.2088

OpenStudy (anonymous):

so then I take .5000-0.2088?

OpenStudy (amistre64):

good, now to adjust this with our zscore ... a good way to think of this is to always start at .5 .5 now notice the sign on you zscore; its a - in this case, so build it in .5 - now we can include the zvalue from the table .5 - .2088 and that should give us the answer they want

OpenStudy (anonymous):

0.2912

OpenStudy (anonymous):

so: P(x<68)=0.2912

OpenStudy (amistre64):

without any computers or calculators to get it more precise, id say thats good

OpenStudy (anonymous):

ok, you want me to type another question? so start a new one or stay on this one?

OpenStudy (amistre64):

start a new one is best, allows others to join in if need be

OpenStudy (anonymous):

ok, im going to type it up.